OpenStudy (anonymous):

which quadratic equation has a vertex of (0,5)

OpenStudy (anonymous):

okay what are the equation

OpenStudy (anonymous):

y=(x-5)^2 y=(x+5)^2 y=x^2-5 y=x^2+5

OpenStudy (dumbcow):

\[y = a(x-h)^{2} +k\] vertex at (h,k)

OpenStudy (anonymous):

so which equation is that

OpenStudy (dumbcow):

thats the general quadratic equation in vertex form. they gave you a vertex --> h=0, k=5 which equation has h=0 and k=5 using the general form as a guide

OpenStudy (anonymous):

y=(x+5)^2 ?

OpenStudy (dumbcow):

almost, you have the right idea but vertex is (h,k) h is x_coordinate h = 0 so it follows that (x-h)^2 is (x-0)^2 = x^2

OpenStudy (anonymous):

y=(x-5)^2 ?

OpenStudy (dumbcow):

ok look at general equation (a = 1) --> (x-h)^2 + k we know h=0 and k =5 from the vertex given (0,5) plug it in -->(x-0)^2 + 5

OpenStudy (anonymous):

so it must be y=x^2+5

OpenStudy (dumbcow):

yes